### v2

# Limitations of Data-Driven Spectral Reconstruction – Optics-Aware Analysis and Mitigation

Abstract.
Hyperspectral imaging empowers machine vision systems with the distinct capability of identifying materials through recording their spectral signatures. Recent efforts in data-driven spectral reconstruction aim at extracting spectral information from RGB images captured by cost-effective RGB cameras, instead of dedicated hardware.
In this paper we systematically analyze the performance of such methods, evaluating both the practical limitations with respect to current datasets and overfitting, as well as fundamental limitations with respect to the nature of the information encoded in the RGB images, and the dependency of this information on the optical system of the camera.
We find that, the current models are not robust under slight variations, e.g., in noise level or compression of the RGB file. Without modeling underrepresented spectral content, existing datasets and the models trained on them are limited in their ability to cope with challenging metameric colors. To mitigate this issue, we propose to exploit the combination of metameric data augmentation and optical lens aberrations to improve the encoding of the metameric information into the RGB image, which paves the road towards higher performing spectral imaging and reconstruction approaches.
Hyperspectral imaging, Spectral reconstruction from RGB, Metamerism, Overfitting, Aberration.
ccs: Computing methodologies Hyperspectral imagingccs: Computing methodologies Computational photography
Figure 1. Spectral image formation models used in the analysis in this work. Top: In the NTIRE spectral recovery challenges, an RGB image is considered as a linear projection from a high-dimensional hyperspectral datacube to a 3D color image. The existence of metamerism results in identical RGB images for different spectra. The neural network trained in this way cannot distinguish their corresponding spectra. Bottom: A possible mitigation to this problem is to include the optical aberrations of the lens in the image formation model. Spectral information is encoded into the aberrated RGB images, enabling the neural network to tell the difference between metamers. In both cases, the RGB image differences are shown on the right (intensity enhanced for better visualization).
1. Introduction
(Hyper)spectral imaging is a method that involves recording the light in a scene in the form of many, relatively narrow, spectral bands, rather than projected into three broadband RGB color channels. Where RGB imaging utilizes the trichromaticity theory of human color vision, spectral imaging provides additional information that can help discriminate between different materials and lighting conditions that are hard to tell apart in RGB images. For example, red stains in a crime scene could be blood, or paint, or a dyed cloth, which cannot be distinguished from their RGB colors. Skin tumors could not be diagnosed from surrounding tissues of the same color. It is difficult to spot and sort out plastic leaves from living plants by their greenish colors. Therefore, spectral imaging has been applied in many fields, including computer graphics (Kim, 2013), machine vision (Fauvel et al., 2012; Rangnekar et al., 2022), healthcare (Lu and Fei, 2014), agriculture (Dale et al., 2013), and environment (Banerjee et al., 2020), to name just a few.
However, conventional hyperspectral cameras require scanning mechanisms (Lodhi et al., 2019; Huang et al., 2022) to acquire the 3D hyperspectral datacube with 2D sensors. To simplify the demanding hardware, extensive efforts have been made in the development of various snapshot hyperspectral cameras (Wagadarikar et al., 2008; Saragadam and Sankaranarayanan, 2019; Jeon et al., 2019).
On the extreme end of these hardware simplification efforts, deep learning methods have emerged in recent years that attempt to solve the problem entirely in software by reconstructing spectral data from RGB images. This has resulted in three CVPR-hosted NTIRE challenges (Arad et al., 2018, 2020, 2022) and various network architectures (Shi et al., 2018; Li et al., 2020; Zhao et al., 2020; Cai et al., 2022b; Zhang et al., 2022). Yet it remains unclear how these methods generalize to unseen data, how they deal with the difficult but important problem of resolving metamerism (Hill, 1999; Ortega et al., 2020), and how they depend on the optical system of both the RGB source and the spectral cameras used to capture the datasets.
At its core, estimating spectral information from RGB colors is an ill-conditioned one-to-many mapping problem. As stated By Pharr et al. (Pharr et al., 2023) (Ch. 4), “any such conversion is inherently ambiguous due to the existence of metamers”. Intuition would therefore suggest that the achievable spectral fidelity of RGB to spectral methods is limited. However, this intuition flies in the face of very high numerical test results reported in recent NTIRE challenges (Arad et al., 2018, 2020, 2022).
One possible explanation for the experimental success of spectral reconstruction from an RGB image is that the networks learn to exploit spatial structure to reconstruct spectral information. However, in practice metameric or near-metameric colors often occur in situations where the spatial structure is also similar, e.g. vein finding (the global geometry is a segment of a forearm, but vein structure is unknown), or biometrics (e.g., distinguishing real faces from masks or images). To illustrate this effect, we show an example from the smaller and older CAVE dataset (Yasuma et al., 2010) fake_and_real_food_ms, where several fake and real objects are presented in two groups, as shown in Fig. 2. Two sample points from the two red peppers (one real and the other fake, but it’s unclear which is which) have the RGB values [86, 21, 10] and [86, 21, 8], respectively. As can be seen, the real spectra of the real and fake red peppers differ substantially in the red part of the spectrum. A pre-trained model (arbitrarily chosen from previous NTIRE challenges trained on the much larger and newer ARAD1K dataset (Arad et al., 2022)) predicts spectra that differ from the ground truth, but more importantly, the predicted spectra for the real and fake pepper are almost identical, illustrating the failure to adequately deal with metamers.
Refer to caption fake_and_real_food_ms spectra
Figure 2. An example scene fake_and_real_food_ms from the CAVE dataset (Yasuma et al., 2010) consists of objects with visually similar colors, but actually different spectra. Left: Color image with highlighted points on the red peppers. Their RGB values are nearly the same. Right: Ground-truth and reconstructed spectra at the corresponding points show their spectral differences. The reconstructed spectra are predicted by the pre-trained MST++ model (Cai et al., 2022b) on the ARAD1K dataset (Arad et al., 2022). The neural network struggles to distinguish either the two spectra from each other, or from their true spectra.
In this paper, we take a systematic and outside-the-box look at all the above aspects. To the best of our knowledge, we are the first to analyze, document, and discuss the inherent shortcomings of this research theme. We highlight realistic conditions under which recent efforts fall short, aiming to constructively instigate, debate, provide insights, and forge a new path regarding the physical phenomena that have been overlooked. By conducting a series of adversarial attacks and thorough analysis, we reveal a number of shortcomings in both current datasets and reconstruction methods. Specifically, we find that:
  • Existing hyperspectral image datasets severely lack in diversity especially with respect to metameric colors but also other factors including nuisance parameters such as noise and compression ratios.
  • State-of-the-art methods suffer from atypical overfitting problems that arise from various factors in the image simulation pipeline, such as noise, RGB data format, and lack of optical aberrations.
  • Optical aberrations in RGB images, while currently ignored by all methods, are actually beneficial rather than harmful to spectral reconstruction if modeled accurately.
  • Crucially, the limitations of the datasets that we document not only affect the RGB to spectral work, but also any other spectral reconstruction and processing that uses the same training data (Jacome et al., 2023; Xu et al., 2024; Arguello et al., 2021). We show how metameric augmentation can be used to at least partially overcome the dataset issues.
Through the evidence in this work, we contribute to a deepened understanding of the limitations of current datasets as well as of underlying sources that result in the limitations of spectral reconstruction accuracy. The results of the interplay between metameric spectra and optical aberrations opens the door for new approaches for spectral recovery down the road.
2. Related Work
2.1. Hyperspectral cameras
Conventional hyperspectral imaging systems require filter wheels, liquid-crystal tunable filters, or mechanical motion (e.g., pushbroom) (Lodhi et al., 2019; Huang et al., 2022) to scan the 3D hyperspectral datacube. To enable snapshot acquisition, coded-aperture snapshot spectral imager (CASSI) (Wagadarikar et al., 2008; Choi et al., 2017b) has been proposed to achieve high spectral accuracy using spectrum-dependent coded patterns. Recent methods also exploit spectrally encoded point spread functions (PSFs) to computationally reconstruct a hyperspectral image (Jeon et al., 2019; Baek et al., 2017; Cao et al., 2011). In general, great efforts have been made to simplify hyperspectral camera hardware by software reconstruction.
2.2. Spectral reconstruction from RGB images
A recent trend to solve the snapshot hyperspectral imaging problem is to exploit hyperspectral data with deep neural networks to reconstruct spectral information from RGB images (Arad and Ben-Shahar, 2016). Owing to the wide availability of RGB cameras, this approach seems to be a promising candidate for hyperspectral imaging if successful. A large number of neural network architectures have been proposed in the past three NTIRE spectral recovery challenges  (Arad et al., 2018, 2020, 2022). Our analysis in this paper focuses on this class of methods to gain insights on their strengths and limitations.
2.3. Dataset bias and data augmentation
Deep neural networks are prone to suffer from data bias (Torralba and Efros, 2011; Fabbrizzi et al., 2022) and overfitting problems (Bejani and Ghatee, 2021). Overfitting can lead to the inability of trained models to generalize in real-world applications (Kornblith et al., 2019). Although overfitting can sometimes be detected by inspecting the training and validation performance over the course of training, it can often be imperceivable in challenging problems. A useful technique to detect overfitting is to use adversarial examples (Werpachowski et al., 2019) generated from the original dataset. On the other hand, it is important to address overfitting when the amount of data is limited. Data augmentation (Shorten and Khoshgoftaar, 2019; Rebuffi et al., 2021) techniques are usually employed to improve the robustness of deep neural networks.
2.4. Metamerism
Metamerism is a physical phenomenon where distinct spectra produce the same color (Akbarinia and Gegenfurtner, 2018) as the high-dimensional spectral space is projected down to three dimensions of a trichromatic vision system (either the human eye or an RGB camera). This phenomenon has been studied in color science (Thornton, 1998; Foster et al., 2006), spectral rendering (Jakob and Hanika, 2019; Weidlich et al., 2021; Van De Ruit and Eisemann, 2023), and hyperspectral imaging (Hill, 1999; Foster and Amano, 2019). In hyperspectral imaging (Cukierski and Foran, 2010; Ortega et al., 2020), it is crucial for many applications to distinguish between metamers or near-metamers (i.e., different spectra that project to similar RGB values). Indeed spectral imaging is usually employed when the RGB color differences between two materials or features are too small to reliably distinguish between them. Therefore, hyperspectral imaging systems require special attention in the system design to acquire accurate spectral signatures (Hill, 2002; Ji et al., 2023).
3. Fundamentals
3.1. Spectral image formation
We denote the hyperspectral image as a matrix ${\mathbf{X}}\in\mathbb{R}^{MN\times K}$, where $M,N$ are the number of pixels, and $K$ is the number of spectral bands. Note that we have stacked the 2D spatial dimensions in rows of ${\mathbf{X}}$. The spectral response function (SRF) of the camera can be expressed as a matrix ${\mathbf{Q}}\in\mathbb{R}^{K\times 3}$. Therefore, the spectrum-to-color projection results in a color image
(1) ${\mathbf{Y}}={\mathbf{X}}{\mathbf{Q}},$
where ${\mathbf{Y}}\in\mathbb{R}^{MN\times 3}$. This is the color formation model in the NTIRE 2022 challenge (Arad et al., 2022). The inverse problem is to recover ${\mathbf{X}}$ from ${\mathbf{Y}}$.
In the past NTIRE challenges (Arad et al., 2018, 2020, 2022), optical aberrations haven’t been included in the image simulation pipeline. However, the optical system of the RGB camera inevitably introduces spectrally-varying blurs to the spectral images, which is modeled as PSFs. This optical process can be described by a linear matrix-vector product in each spectral band followed by a sum over the spectral dimension. The spectral images through the optical system are ${\mathbf{W}}=\mathrm{diag}\left({\mathbf{A}}{\mathbf{X}}\right)$, where ${\mathbf{A}}\in\mathbb{R}^{KMN\times MN}$ is a block matrix that stacks the spectral PSF matrices vertically, and $\mathrm{diag}(\cdot)$ extracts the diagonal blocks. The final RGB image is then
(2) ${\mathbf{Z}}=\mathrm{diag}\left({\mathbf{A}}{\mathbf{X}}\right){\mathbf{Q}},$
where ${\mathbf{Z}}\in\mathbb{R}^{MN\times 3}$. With the optical image formation model accounted for, the inverse problem is to recover ${\mathbf{X}}$ from ${\mathbf{Z}}$. See Appendix for the detailed image formation models.
The two spectral image formation models are illustrated in Fig. 1. Without considering optical aberrations (Eq. (1)), the neural network struggles to reconstruct the real hyperspectral images in the presence of metamers. The spectrally-varying PSFs (Eq. (2)) are helpful to mitigate this issue since the aberrated RGB images from metamers are different. In the following sections, we will discuss the limitations of existing data-driven spectral reconstruction based on these two image formation models in detail.
3.2. Hyperspectral datasets and data diversity
Compared to very large color (RGB) image datasets (e.g., ImageNet (Deng et al., 2009), DIV2K (Agustsson and Timofte, 2017)), hyperspectral datasets are far smaller in size, primarily limited by the unavailability of high-quality hyperspectral cameras and the difficulty in acquiring outside the lab with moving target scenes. The largest dataset so far is ARAD1K used in the NTIRE 2022 challenge (Arad et al., 2022). In addition, we also include the CAVE (Yasuma et al., 2010), ICVL (Arad and Ben-Shahar, 2016), and KAUST (Li et al., 2021) datasets that share the same spectral range to extend our experiments. The datasets are summarized in Table 1. Although other datasets (Chakrabarti and Zickler, 2011; Choi et al., 2017a; Monno et al., 2015) exist, they cover slightly different spectral bands, making direct comparisons difficult. We therefore restrict our analysis to the datasets listed in the table.
Table 1. Basic information of four existing hyperspectral datasets.
Dataset Spectra (nm) Resolution (x, y, $\lambda$) Amount Device Scene
CAVE (Yasuma et al., 2010) 400:10:700 512 $\times$ 512 $\times$ 31 32 monochrome sensor + tunable filters lab setup
ICVL (Arad and Ben-Shahar, 2016) 400:10:700 1392 $\times$ 1300 $\times$ 31 201 HS camera (Specim PS Kappa DX4) outdoor
KAUST (Li et al., 2021) 400:10:700 512 $\times$ 512 $\times$ 31 409 HS camera (Specim IQ) outdoor
ARAD1K (Arad et al., 2022) 400:10:700 482 $\times$ 512 $\times$ 31 1000 HS camera (Specim IQ) outdoor
The difficulties in the data capture not only affect the size but also the diversity of the datasets. In particular, effects like metamerism, which are comparatively rare in everyday environments (Foster et al., 2006; Prasad and Wenhe, 2015), yet crucial for many actual applications of spectral imaging, are under-represented in the datasets. While the CAVE dataset (Yasuma et al., 2010) contains some fake-and-real pairs of objects (e.g., Fig. 2) to account for metamerism, the total amount of such data is still very low. We analyze the general data diversity issue in Section 4 and the specific case of metamerism in Section 5.
3.3. Modeling metamerism
Since there are not enough examples of metamerism in existing datasets, their effects in spectral reconstruction went unnoticed in prior works. On the one hand, we need adversarial examples to reveal the unexplored problems. On the other hand, we want to investigate how they can complement the current datasets. Therefore, we propose a new form of data augmentation in our experiments. Metameric augmentation starts with existing spectral images and creates a new, different spectral image that however maps to the same RGB image (given a specific set of RGB spectral response functions). In Section 5, we first use metameric augmentation as an adversarial example to reveal the previously omitted effects of metamerism on the performance discrepancy. In Section 6, we show metameric augmentation is beneficial to mitigate the raised problems along with an aberration-aware training strategy.
Note that data augmentation has proven to be effective in deep learning to mitigate data shortage. Color image augmentation techniques have been focusing mainly on geometric transformations and intensity adjustment. Although these techniques have been employed in prior methods, they only augment the spatial dimensions in hyperspectral images. To the best of our knowledge, metameric augmentation beyond RGB colors, accounting for the physical phenomenon of metamerism, has not been adopted before in spectral reconstruction problems. Metameric augmentation can greatly enrich the spectral content in existing datasets to mitigate the lack of diversity.
Interestingly, metamer generation from existing spectra has been studied in color science and spectral rendering to accurately model the scenes using various methods, e.g., metameric black (van Trigt, 1994; Finlayson and Morovic, 2005; Alsam and Lenz, 2007) and spectral uplifting (Jakob and Hanika, 2019; Van De Ruit and Eisemann, 2023; Belcour et al., 2023). In this work, to support our analysis, we adopt the metameric black approach to generate metamers, whereas we note that other metamer generation methods can also be employed for the same purpose.
A spectrum ${\mathbf{S}}$ can be projected onto two orthogonal subspaces, one for the fundamental metamer ${\mathbf{S}}^{*}$, and the other for metameric black ${\mathbf{B}}$ (Cohen and Kappauf, 1982; Viénot and Brettel, 2014). A new metameric spectrum ${\mathbf{S}}^{\prime}$ is then
(3) ${\mathbf{S}}^{\prime}={\mathbf{S}}^{*}+\alpha{\mathbf{B}},$
where ${\mathbf{S}}^{*}={\mathbf{Q}}\left({\mathbf{Q}}^{T}{\mathbf{Q}}\right)^{-1}{\mathbf{Q}}^{T}{\mathbf{S}}$ and ${\mathbf{B}}={\mathbf{S}}-{\mathbf{S}}^{*}$. Since adding metameric black does not alter the RGB color, we can vary the coefficient $\alpha$ to generate different spectra that are all metamers. To avoid negative spectral radiance, we clip the negative values in the generated spectra and re-calculate the RGB colors for the affected pixels. See Appendix for the analysis on the effects of clipping to non-negative values while generating metamer data.
3.4. Performance evaluation metrics
Consider a hyperspectral image ${\mathbf{X}}_{k,i,j}$ and its estimate $\hat{{\mathbf{X}}}_{k,i,j}$, where $k$ is the spectral index, and $i,j$ are spatial indices. The reconstruction quality can be evaluated in various ways. The NTIRE 2022 spectral reconstruction challenge (Arad et al., 2022) adopts two numerical metrics, Mean Relative Absolute Error (MRAE),
(4) ${\mathrm{\mathbf{MRAE}}}=\frac{1}{KMN}\sum_{k,i,j}\left|\hat{{\mathbf{X}}}_{k,i,j}-{\mathbf{X}}_{k,i,j}\right|/{\mathbf{X}}_{k,i,j},$
and Root Mean Square Error (RMSE)
(6) ${\mathrm{\mathbf{RMSE}}}=\sqrt{\frac{1}{KMN}\sum_{k,i,j}\left(\hat{{\mathbf{X}}}_{k,i,j}-{\mathbf{X}}_{k,i,j}\right)^{2}}.$
Another metric widely used in hyperspectral imaging is the Spectral Angle Mapper (SAM) (Van der Meer, 2006; Kuching, 2007; Park et al., 2007), although it has not yet found its way into the relevant computer vision literature. SAM emphasizes the spectral accuracy compared to the previous metrics, which are more forgiving of large errors in individual spectral channels:
(7) ${\mathrm{\mathbf{SAM}}}=\frac{1}{MN}\sum_{i,j}\cos^{-1}\left(\frac{\sum_{k}\hat{{\mathbf{X}}}_{k,i,j}{\mathbf{X}}_{k,i,j}}{\sqrt{\sum_{k}\hat{{\mathbf{X}}}_{k,i,j}^{2}\sum_{k}{\mathbf{X}}_{k,i,j}^{2}}}\right).$
Finally, we also inspect the spatial quality in individual spectral channels, and calculate the spectrally averaged Peak Signal-to-Noise Ratio (PSNR),
(8) ${\mathrm{\mathbf{PSNR}}}=\frac{1}{K}\sum_{k}20\log_{10}\left(\frac{\textrm{MAX}}{\sqrt{\textrm{MSE}_{k}}}\right),$
where MAX is the maximum possible value, and $\textrm{MSE}_{k}$ is the mean squared error in the $k$-th spectral band. This metric complements RMSE to account for performance variation in individual spectral bands.
3.5. Training details
In our study, we conduct all the experiments across various datasets and network architectures. Following the methodologies proposed by the latest champion network MST++ (Cai et al., 2022b), we employ their patch-wise training approach (patches of 128$\times$128 pixels).
3.5.1. Data Preparation
Throughout the experiments in this work, we follow the same data format (Matlab-compatible mat files) for hyperspectral images in the ARAD1K dataset (Arad et al., 2022). To be consistent, we also convert the raw hyperspectral datacubes in the CAVE (Yasuma et al., 2010), ICVL (Arad and Ben-Shahar, 2016), and KAUST (Li et al., 2021) datasets to this format. The data values are normalized by their respective bit-depths such that the data range is [0.0, 1.0]. The training and validation sets in ARAD1K are kept the same as offered in the NTIRE 2022 spectral recovery challenge (Arad et al., 2022), i.e., 900 files for training, and 50 for validation. We split the CAVE, ICVL, and KAUST datasets by 90% for training, and 10% for validation. Following the training strategy in MST++ (Cai et al., 2022b), we keep the training and validation lists fixed.
3.5.2. Metamer Generation
We adopt the metameric black method (van Trigt, 1994; Finlayson and Morovic, 2005; Alsam and Lenz, 2007) to generate metamers from the original hyperspectral data. By varying the coefficient of the metameric black term, we could generate metamers that project to the same RGB color. Note that the original datacube corresponds to $\alpha=1$, and the fundamental metamer corresponds to $\alpha=0$. For the experiments with fixed metamers, we use the fundamental metamers to complement the original standard data. This is sufficient to demonstrate our findings. Other arbitrary values would result in the same conclusions. A more aggressive setting is to vary $\alpha$ as a variable to account for the infinite possible metamers in a more realistic situation.
3.5.3. Training and Validation Procedure
Following the training strategies in MST++ (Cai et al., 2022b), we sub-sample the hyperspectral datacubes and the corresponding RGB images into overlapping patches of 128$\times$128. Spatial augmentations, such as random rotation, vertical flipping, and horizontal flipping are randomly applied to the training patches. In the validation step, we calculate the evaluation metrics (MRAE, RMSE, PSNR, and SAM) on the full spatial resolution for the ARAD1K (482$\times$512), CAVE (512$\times$512), and KAUST (512$\times$512) datasets. Note that this is different from MST++ (Cai et al., 2022b) where only the central 256$\times$256 regions are evaluated. The ICVL dataset has a very large spatial resolution (1300$\times$1392). To be consistent with other datasets, we evaluate ICVL only in the central 512$\times$512 regions.
Similar as MST++, in each epoch, we train the networks for 1000 iterations, with a total number of 300 epochs. All the reported results are evaluated at the end of the training epochs, i.e., 300k iterations. We find that the training iterations are sufficient to achieve convergence in all our proposed experiments. Hyperparameters, such as learning rate and batch size, are tuned to achieve the best performance for each network on each dataset.
All the experiments are conducted on an NVIDIA A100 GPU (80 GB memory).
4. Finding 1: Atypical Overfitting
Although it is well-known that deep neural networks may suffer from overfitting problems, we find that the overfitting behavior in spectral reconstruction is atypical and difficult to notice with standard evaluations. Here we introduce minimalist changes to the ARAD1K dataset used in NTIRE 2022 challenge (Arad et al., 2022) in three experiments to demonstrate it. We evaluate 11 top-performant neural network architectures, namely MST++ (Cai et al., 2022b), MST-L (Cai et al., 2022a), MPRNet (Zamir et al., 2021), Restormer (Zamir et al., 2022), MIRNet (Zamir et al., 2020), HINet (Chen et al., 2021), HDNet (Hu et al., 2022), AWAN (Li et al., 2020), EDSR (Lim et al., 2017), HRNet (Zhao et al., 2020), and HSCNN+ (Shi et al., 2018).
4.1. Training with less data
First, we make a simple change to the training of the participating networks in the NTIRE 2022 challenge (Arad et al., 2022). While keeping all the training settings intact, we randomly choose only 50% or 20% of the original training data, respectively, to train the candidate networks and validate the performance on the original validation data. We illustrate the validation curves for MST++ (Cai et al., 2022b) in Fig. 3. See the Supplemental Material for the results of other networks. We summarize the results for 100% and 50% training data in Table 2 for all networks.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3. Validation performance for MST++ (Cai et al., 2022b) with 100%, 50%, and 20% of the original training data on ARAD1K (Arad et al., 2022).
Although the performance with less training data deviates mildly in MRAE, RMSE, and PSNR, the spectral accuracy SAM (highlighted in bold in Table 2) is surprisingly less affected. In particular, some networks (e.g., MST++ (Cai et al., 2022b), MIRNet (Zamir et al., 2020)) achieve exactly the same SAM scores. MST-L (Cai et al., 2022a) (50%) even improves SAM slightly, placing itself the best among all.
Table 2. Performance comparison with 100% and 50% of training data for different methods on the original ARAD1K dataset.
[…]
MST++ 100% 0.182 0.033 33.0 0.086
50% 0.209 0.033 32.7 0.086
MST-L 100% 0.184 0.031 33.5 0.084
50% 0.253 0.042 30.6 0.080
MPRNet 100% 0.212 0.034 32.5 0.084
50% 0.293 0.039 31.3 0.091
Restormer 100% 0.204 0.033 33.2 0.083
50% 0.304 0.041 31.4 0.092
MIRNet 100% 0.186 0.030 33.7 0.082
50% 0.214 0.033 32.6 0.082
HINet 100% 0.234 0.036 32.3 0.085
50% 0.267 0.041 30.7 0.090
HDNet 100% 0.223 0.038 31.2 0.095
50% 0.296 0.047 28.9 0.097
AWAN 100% 0.213 0.034 32.2 0.091
50% 0.273 0.042 30.5 0.095
EDSR 100% 0.358 0.052 27.3 0.095
50% 0.430 0.059 26.1 0.093
HRNet 100% 0.388 0.057 26.4 0.094
50% 0.413 0.065 25.5 0.096
HSCNN+ 100% 0.428 0.066 25.4 0.098
50% 0.462 0.068 25.0 0.101
We therefore paradoxically find that despite the small size of hyperspectral datasets, the data already seems to be redundant. This serves as a first indication that the diversity of the datasets is severely lacking. We analyze this effect in more depth in the following experiments.
4.2. Validation with unseen data
Table 3. Evaluation of the pre-trained MST++ model on synthesized validation data from the ARAD1K dataset with different noise levels, compression quality, and realistic optical aberrations.
Data property MRAE $\downarrow$ RMSE $\downarrow$ PSNR $\uparrow$ SAM $\downarrow$
Data source Noise (npe) RGB format Aberration
1 NTIRE 2022 unknown jpg (Q unknown) None 0.170 0.029 33.8 0.084
2 Synthesized 0 jpg (Q = 65) None 0.460 0.049 29.2 0.094
3 0 png (lossless) None 0.362 0.057 28.7 0.087
4 1000 png (lossless) CA* 0.312 0.055 28.4 0.118
To further scrutinize the underlying issue, we validate existing pre-trained models with “unseen” data synthesized from the original dataset used in the NTIRE challenge. The challenge organizers state that
the exact noise parameters and JPEG compression level used to generate RGB images for the challenge was kept confidential
(Arad et al., 2022). Only the spectrum-to-color projection was considered, and no aberrations of the optical system were simulated.
In our experiments, we generate new RGB images using the same methodology and calibration data, but different noise and compression settings. Specifically, we use the SRF data for a Basler ace 2 camera (model A2a5320-23ucBAS) known to the networks, and simulate Poisson noise at varying noise levels by controlling the number of photon electrons (npe). We adopt the same rudimentary in-camera image signal processing pipeline. As an illustrative example, we use MST++ (Cai et al., 2022b) in Table LABEL:tab:unseen, Row 1 as a reference for comparison; results for other networks can be found in the Supplemental Material.
First, as a baseline, we consider a noiseless (npe = 0) and aberration-free case with moderate JPEG compression quality (Q = 65), shown in Table LABEL:tab:unseen, Row 2. The results show significant drops in all the performance metrics. Note that the only differences here compared to the challenge dataset are the noise level and compression quality – the base images are identical! This indicates that the network overfits both the noise and JPEG compression parameters.
Second, in Row 3, we generate noiseless RGB images, but in lossless PNG format, as opposed to the JPEG (Q = 65) in Row 2. Note that JPEG compression is not necessary for the core inverse problem in hyperspectral imaging, since raw data could be readily obtained from the sensors. This results in paradoxical reconstruction performance. MRAE and SAM improve compared to Row 2 (but are still worse than Row 1), while RMSE and PSNR deteriorate further. Considering this only eliminates image compression, and the networks were trained on MRAE (Cai et al., 2022b), we can confirm that the network indeed overfits the specific unknown JPEG compression used in the challenge (Arad et al., 2022).
Third, we consider a more realistic imaging scenario, in which we eliminate the impact of unnecessary compression by employing the lossless PNG format to save the RGB images (equivalent to using raw camera data). We adopt moderate noise levels (npe = 1000) and realistic optical aberrations from a recent double Gauss lens patent (Ichimura, 2021) to mimic a real photographic camera. We can observe a further performance drop in Row 4, which provides additional evidence that the network overfits the unknown parameters in the image simulation pipeline (Arad et al., 2022). When used under realistic imaging conditions, the performance degrades significantly.
4.3. Cross-dataset validation
In addition, we inspect the effects of different datasets (cf. Table 1) on the performance. We train the MST++ network on the four datasets with the same image simulation parameters. To eliminate the impact of other factors, we choose the ideal noiseless and aberration-free condition without compression. In the validation, we use our trained model on ARAD1K dataset to validate on the other three datasets, respectively. In Table 4, we compare the performance with the models both trained and validated on the original datasets. Results for other networks can be found in the Supplement. They all illustrate the same difficulties in generalization.
Table 4. Cross-dataset validation example using MST++ (Cai et al., 2022b).
[…]
CAVE CAVE 0.237 0.034 31.9 0.194
ARAD1K CAVE 1.626 0.074 24.4 0.376
ICVL ICVL 0.079 0.019 38.3 0.024
ARAD1K ICVL 0.627 0.091 22.0 0.110
KAUST KAUST 0.069 0.013 44.4 0.061
ARAD1K KAUST 1.042 0.100 22.0 0.370
Even though the imaging conditions are the same and ideal, the network trained on one dataset experiences significant performance drop in all metrics when validated on other datasets. This indicates that the contents of the datasets, as well as the acquisition devices used to capture the datasets, play important roles.
We also point out that the CAVE dataset (Yasuma et al., 2010), although smaller and older than the others, is more difficult to train for better performance. This is probably due to the fact that CAVE consists of several challenging scenes of real and fake objects that appear in similar colors, but other datasets comprise less aggressive natural scenes.
4.4. Discussion
The experiments conducted in this section clearly highlight several shortcomings with respect to the existing datasets: (1) they lack diversity in nuisance parameters such as noise and compression ratios, and (2) they lack scene diversity. While we show in this document the results for the largest available dataset (ARAD1K), the Supplement shows consistent results for all the datasets. Both of these aspects result in over-fitting and prevent the networks from learning the general spectral image restoration task. Next we specifically analyze the effect of metamerism; the analysis of the impact of optical aberrations will be deepened in Section 6.
Refer to caption standard metamer 420 nm 500 nm 550 nm 580 nm 660 nm GT std recon std GT metamer recon metamer
Figure 4. Validation with metamers for MST++ (Cai et al., 2022b). An example Scene ARAD_1K_0944 is shown to visualize the standard and metamer datacubes. Top left: the standard and metamer data result in similar color images. Bottom left: ground-truth and reconstructed spectra from two labeled points. Right: ground-truth and reconstructed spectral images in 420 nm, 500 nm, 550 nm, 580 nm, and 660 nm.
5. Finding 2: Metameric Failure
In this section, we inspect the performance of existing methods using metamer as an adversary to validate as well as re-train the neural networks for performance analysis.
5.1. Validation with metamers
We generate metamer datacubes (metamer data) from the original ARAD1K dataset (standard data) using the metameric black method (Finlayson and Morovic, 2005). For this set of experiments, we fix the coefficient $\alpha=0$ in Eq. (3). We choose a realistic imaging condition as used in Table LABEL:tab:unseen, Row 4, and keep it the same for both cases. The validation results on the ARAD1K dataset for existing pre-trained networks are summarized in Table 5. We also visualize the reconstructed spectral images in five arbitrary bands (420 nm, 500 nm, 550 nm, 580 nm, and 660 nm) and spectra of two points in Fig. 4 for Scene ARAD_1K_0944 from the validation set.
Table 5. Validation performance for different pre-trained models on standard (std) data and metamer (met) adversary synthesized from the ARAD1K dataset (Arad et al., 2022).
[…]
MST++ std 0.312 0.055 33.8 0.084
met 52.839 0.091 26.0 0.580
MST-L std 0.327 0.055 28.0 0.118
met 51.321 0.090 25.9 0.579
MPRNet std 0.661 0.066 26.1 0.125
met 145.981 0.122 23.3 0.547
Restormer std 0.510 0.066 25.5 0.126
met 79.705 0.116 23.4 0.567
MIRNet std 0.404 0.077 24.8 0.124
met 38.252 0.089 24.8 0.570
HINet std 0.450 0.063 26.5 0.120
met 67.148 0.096 24.8 0.552
HDNet std 0.450 0.082 23.9 0.126
met 34.429 0.095 23.8 0.570
AWAN std 0.424 0.080 24.6 0.119
met 39.854 0.095 24.4 0.558
EDSR std 0.421 0.066 25.5 0.132
met 49.435 0.100 23.8 0.564
HRNet std 0.514 0.078 23.9 0.128
met 43.726 0.112 22.7 0.560
HSCNN+ std 0.508 0.075 24.4 0.148
met 42.274 0.098 23.1 0.556
From the numerical results in Table 5, it is apparent that all the existing methods experience catastrophic performance drop in terms of MRAE and SAM in the presence of metamers, which we call metameric failure. The MRAE (cf. Eq. (4)) may yield large values when big errors occur for dark ground-truth pixels (see the exemplary spectra in Fig. 4). The SAM values become large when the spectra are essentially dissimilar with each other. RMSE and PSNR do not capture the spectral differences as well, since they average out differences in the spatial and spectral dimensions.
The visual results in Fig. 4 show that the reconstruction results are very close to each other for both standard and metamer data, because the input RGB images are quite similar. However, distinct differences exist in the scene for certain spectral bands, e.g., the intensities of the yellow and green parts of the slide (blue box) in 500 nm band vary in the standard data, but remain identical in the metamer data. The reconstructions fail to reflect this important difference. All spectral images are displayed on the same global intensity scale, so the brightness differences (green box) in corresponding images reflect the reconstruction artifacts for the metamer data.
5.2. Training with metamers
The pre-trained models were not explicitly trained to cope with metamers. This raises the question whether it is possible to improve the performance by training the networks with metamer data.
As a first step, we use both the standard and metamer data ($\alpha=0$) generated from the ARAD1K dataset to train various networks. To eliminate the impact of other factors, we simulate the RGB images in a noiseless, aberration-free condition, and without compression.
However, it is not sufficient to consider only a pair of standard and fixed metamer data. In reality, there are infinite metamers that project to the same color. As a second variant, we train the neural networks with random metamers generated on-the-fly as a spectral augmentation to enhance the spectral content of existing datasets. We vary the coefficient for the metameric black by setting $\alpha$ as a uniformly distributed random number in the range $[-1,2]$. During validation, we use both the standard validation data and their corresponding metamer data with fixed $\alpha=0$, which doubles the amount of the original validation data.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 5. Training MST++ with metamers. It fails to combat fixed metamers and on-the-fly metamers, in particular on the spectral accuracy SAM.
As an example, we train MST++ and evaluate its validation performance over the training process. In Fig. 5, we show that it is no longer a good choice to use MRAE as the loss function (Cai et al., 2022b) and evaluation metric (Arad et al., 2022), because it is completely overwhelmed by metamers. Instead, we find that L1 loss is a more stable loss function, so we train the network with L1 loss for three cases, no metamer (easy), fixed metamer (medium), and on-the-fly metamer (difficult). Nevertheless, we can see in Fig. 5 that the network fails in particular for the spectral accuracy SAM.
Table 6. Performance comparison for training various networks with fixed and on-the-fly metamers.
[…]
MST++ no 0.270 0.041 31.6 0.079
fixed 9.912 0.056 31.6 0.247
on-the-fly 14.224 0.091 26.1 0.284
MST-L no 0.269 0.040 32.2 0.081
fixed 11.398 0.061 30.2 0.289
on-the-fly 12.155 0.061 27.0 0.258
MPRNet no 0.346 0.051 29.9 0.076
fixed 7.359 0.059 30.6 0.224
on-the-fly 13.492 0.087 26.6 0.264
Restormer no 0.286 0.041 31.5 0.068
fixed 8.129 0.059 31.0 0.242
on-the-fly 10.186 0.089 26.7 0.264
MIRNet no 0.258 0.040 32.2 0.083
fixed 9.555 0.061 30.5 0.289
on-the-fly 13.205 0.088 26.4 0.290
HINet no 0.315 0.056 28.0 0.081
fixed 8.322 0.068 29.6 0.296
on-the-fly 14.238 0.090 25.5 0.288
HDNet no 0.287 0.045 30.2 0.087
fixed 8.884 0.064 30.2 0.296
on-the-fly 18.270 0.087 26.4 0.299
AWAN no 0.240 0.039 32.0 0.073
fixed 8.789 0.068 29.6 0.294
on-the-fly 14.406 0.090 26.0 0.264
EDSR no 0.415 0.061 26.1 0.084
fixed 9.717 0.073 26.6 0.297
on-the-fly 11.723 0.101 23.2 0.290
HRNet no 0.430 0.065 25.6 0.085
fixed 8.142 0.076 26.0 0.299
on-the-fly 13.098 0.103 23.0 0.294
HSCNN+ no 0.516 0.077 24.1 0.082
fixed 9.362 0.085 24.8 0.297
on-the-fly 13.311 0.102 22.9 0.286
We also train all other candidate networks with fixed and on-the-fly metamers. The results are summarized in Table 6. Again, the same performance drop applies to all networks. Finally, we show the results of the top-performing network, MST++ on the CAVE, ICVL, and KAUST datasets in Table 7. (See Supplemental Material for more results). As before, the performance drops similarly in the presence of metamers.
5.3. Discussion
The experiments conducted in this section clearly highlight the difficulties that the data-driven spectral recovery methods face with metamers: (1) lack of sufficient metameric data in current datasets, (2) training with metamers alone cannot mitigate the issue, and (3) spectral estimation from RGB data is indeed limited in the presence of metamers.
The limitations of spectral estimation from RGB data are ultimately not overly surprising – after all the projection from the high dimensional spectral space to RGB invariably destroys scene information that can be difficult to recover. Spatial context from underrepresented data does not contribute to the spectral estimation, because such information remains the same for metamers. However, our experiments show that this is indeed an issue faced by the state-of-the-art methods, which so far went unnoticed due to the under-representation of metamers in the datasets. This shortcoming will also affect other uses of the same datasets, for example in the training of reconstruction methods for spectral computational cameras (Jeon et al., 2019; Baek et al., 2017; Cao et al., 2011). Metameric adversary helps to identify this overlooked issue, whereas metameric augmentation has not yet found its way for contribution. We will explore it further in the next section.
Table 7. Training with metamers for MST++ on the CAVE (Yasuma et al., 2010), ICVL (Arad and Ben-Shahar, 2016), and KAUST (Li et al., 2021) datasets.
[…]
CAVE no 1.014 0.038 29.9 0.192
fixed 38.26 0.053 29.6 0.229
on-the-fly 226.0 0.078 25.2 0.451
ICVL no 0.067 0.016 40.1 0.027
fixed 1.454 0.041 34.8 0.229
on-the-fly 2.615 0.087 24.3 0.268
KAUST no 0.082 0.016 43.2 0.076
fixed 2.033 0.022 39.0 0.217
on-the-fly 1.874 0.032 33.7 0.245
6. Finding 3: The Aberration Advantage
As shown so far, the existing methods have difficulties distinguishing metamers in the ideal noiseless and aberration-free condition. In this section, we analyze what effect (if any) optical aberrations have on this situation, i.e., aberration-aware training (Yang et al., 2023). To this end we train the networks in a realistic imaging condition with moderate noise level (npe = 1000), lossless PNG format, and aberrations from the same double Gauss lens as before (Ichimura, 2021). In short, we simulate, through spectral ray tracing, the effect that an imperfect (i.e., aberrated) optical system has on the RGB image measured when observing a specific spectral scene. The details of this simulation can be found in the Supplemental Material.
In Fig. 6, we show an example with MST++ for the validation on SAM in two situations, one with fixed metamers, and the other with on-the-fly metamers. In each experiment, we compare the SAM difference with and without aberrations. As a reference, we also show the standard validation without metamers. As we can see, the realistic optical aberrations of the lens actually improve the spectral estimation in the presence of metamers as long as the aberrations are modeled in the training. With chromatic aberrations, the network can already distinguish fixed metamer pairs, achieving similar accuracy as the standard case. In the more aggressive case of on-the-fly metamers, chromatic aberrations also improve the spectral accuracy, compared with its no-aberration counterpart. Again, this aberration advantage holds for all datasets (Table 8). See Supplemental Material for details.
Refer to caption Refer to caption
Figure 6. Chromatic aberrations improve spectral accuracy. Left: fixed metamers. Right: on-the-fly metamers.
Table 8. SAM metrics for MST++ on CAVE (Yasuma et al., 2010), ICVL (Arad and Ben-Shahar, 2016), and KAUST (Li et al., 2021).
Dataset Fixed metamers On-the-fly metamers
no aberration aberration no aberration aberration
CAVE 0.251 0.135 0.380 0.167
ICVL 0.028 0.077 0.240 0.085
KAUST 0.212 0.113 0.221 0.113
6.1. Discussion
To understand why optical aberrations help improve the reconstruction, consider the simulated images in Fig. 7. The left and middle images are simulations of RGB images for metameric scene pairs, with the difference image on the right. The different spectra of the two scenes are affected differently by the optical aberrations, and therefore, although the scenes are metamers of each other, the RGB images are in fact different. In effect, the optical aberrations have encoded spectral information into the RGB image, which the networks can learn to distinguish, lending credibility to PSF engineering methods for hyperspectral encoding (Jeon et al., 2019; Baek et al., 2017; Cao et al., 2011).
standard metamer color difference
[…]
Figure 7. Chromatic aberration induced informative color differences (right) as spectral cues for metamer pairs (left and middle).
There are, however, also detrimental consequences of this effect, namely (1) “baked-in” aberrations in spectral datasets and (2) limited generalization to arbitrary RGB image sources.
Regarding (1), it is worth noting that even expensive spectral cameras typically do not have diffraction limited optics, i.e., their optical systems are aberrated. In order to avoid overfitting to these aberrations, future datasets should ideally be collected with a large variety of different spectral cameras.
With respect to (2) we note that different RGB cameras produce very different aberrations, and often the aberrations also vary significantly across the image plane. Without further improvements on the reconstruction methods, this will result in a large device dependency and poor generalization of the methods.
7. Conclusion
In this work, we have comprehensively analyzed a category of data-driven spectral reconstruction methods from RGB images by reviewing the problem fundamentally from dataset bias to physical image formation, and to reconstruction networks. From an optics-aware perspective, we leverage both metamerism and optical aberrations to reassess existing methodologies.
The major findings of our study are that (1) the limitations of current datasets lead to overfitting to both nuisance parameters (noise, compression), as well as limited scene content. (2) Metamerism in particular presents a challenge both in terms of under-representation in the datasets, and in terms of fundamental limits of spectral reconstruction from RGB input. (3) Metameric augmentation along with the targeted use of optical aberrations enables resolving the metamer issue.
In particular the last point lends additional credibility to computational camera approaches for spectral imaging, including both PSF engineering approaches (Jeon et al., 2019; Baek et al., 2017; Cao et al., 2011) and compressed sensing methods like CASSI (Wagadarikar et al., 2008; Choi et al., 2017b). However, learned reconstruction methods for these approaches will suffer from the same dataset issues as the methods analyzed in this paper, making the collection of large-scale, diverse spectral image data a matter of urgency.
In order to realize the dream of spectral estimation from arbitrary RGB sources it is necessary to diversify the spectral contents in hyperspectral image datasets and embrace versatile spectral data augmentation methods to fully enable the power of networks in adaptation to whole families of aberrations.
Acknowledgements.
This work was supported by the King Abdullah University of Science and Technology Individual Baseline Funding. Seung-Hwan Baek was supported by Korea NRF (RS-2023-00211658) and Samsung Electronics.
Appendix
Appendix A Image Formation Model
Mathematically, the physical image formation of a color image from the spectral radiance can be expressed by
[…]
where $f\left(x,y,\lambda\right)$ is the spectral image, $h\left(x,y,\lambda\right)$ is the spectral point spread function (PSF) of the optical system, $q_{c}\left(\lambda\right)$ is the spectral response function (SRF) of the sensor, and $g_{c}\left(x,y\right)$ is the color image in color channel $c\in[R,G,B]$.
Let us denote the hyperspectral image as a matrix ${\mathbf{X}}\in\mathbb{R}^{MN\times K}$, where $M,N$ are the number of pixels in spatial dimensions, and $K$ is the number of spectral bands in spectral dimension. Note that we have stacked the 2D spectral images in rows of ${\mathbf{X}}$. Explicitly, we have
(10) ${\mathbf{X}}=\left[{\mathbf{x}}_{1},{\mathbf{x}}_{2},\dots,{\mathbf{x}}_{K}\right],$
where each column ${\mathbf{x}}_{k}\in\mathbb{R}^{MN\times 1}$ is a vector for the spectral image in spectral channel $k$. The SRF of the sensor is a matrix ${\mathbf{Q}}\in\mathbb{R}^{K\times 3}$, i.e.,
(11) ${\mathbf{Q}}=\left[{\mathbf{q}}_{1},{\mathbf{q}}_{2},{\mathbf{q}}_{3}\right]=\left[\begin{matrix}q_{11}&q_{21}&q_{31}\\ q_{21}&q_{22}&q_{32}\\ \vdots&\ddots&\vdots\\ q_{K1}&q_{K2}&q_{K3}\\ \end{matrix}\right],$
where each column ${\mathbf{q}}_{c}\in\mathbb{R}^{K\times 1}$. Therefore, the spectrum-to-color projection results in a color image
(12) ${\mathbf{Y}}={\mathbf{X}}{\mathbf{Q}},$
where ${\mathbf{Y}}\in\mathbb{R}^{MN\times 3}$ with three columns
(13) ${\mathbf{Y}}=\left[{\mathbf{y}}_{1},{\mathbf{y}}_{2},{\mathbf{y}}_{3}\right],$
and each column ${\mathbf{y}}_{c}\in\mathbb{R}^{MN\times 1}$ is a vector for the image in color channel $c\in[R,G,B]$.
When considering the spectral PSFs in each spectral channel, the optically blurred image can be expressed by
(14) ${\mathbf{w}}_{k}={\mathbf{A}}_{k}{\mathbf{x}}_{k},\quad k\in[1,2,\dots,K],$
where ${\mathbf{A}}_{k}\in\mathbb{R}^{MN\times MN}$ is a matrix that represents the spectral PSF in channel $k$. Similar as ${\mathbf{X}}$, we concatenate ${\mathbf{w}}_{k}$ horizontally to obtain the spectral images through the optical system as
(15) ${\mathbf{W}}$ $=\left[{\mathbf{w}}_{1},{\mathbf{w}}_{2},\dots,{\mathbf{w}}_{K}\right]$
$=\left[{\mathbf{A}}_{1}{\mathbf{x}}_{1},{\mathbf{A}}_{2}{\mathbf{x}}_{2},\dots,{\mathbf{A}}_{K}{\mathbf{x}}_{K}\right].$
We define a block matrix
(16) ${\mathbf{A}}=\left[\begin{matrix}{\mathbf{A}}_{1}\\ {\mathbf{A}}_{2}\\ \dots\\ {\mathbf{A}}_{K}\\ \end{matrix}\right],$
which stacks the matrices ${\mathbf{A}}_{k}$ vertically, and ${\mathbf{A}}\in\mathbb{R}^{KMN\times MN}$. Therefore, we have
(17) ${\mathbf{W}}=\mathrm{diag}\left({\mathbf{A}}{\mathbf{X}}\right),$
where $\mathrm{diag}\left(\cdot\right)$ extracts the $K$ diagonal blocks and concatenate them horizontally,
(18) ${\mathbf{A}}{\mathbf{X}}$ $=\left[\begin{matrix}{\mathbf{A}}_{1}\\ {\mathbf{A}}_{2}\\ \dots\\ {\mathbf{A}}_{K}\\ \end{matrix}\right]\left[{\mathbf{x}}_{1},{\mathbf{x}}_{2},\dots,{\mathbf{x}}_{K}\right]$
$=\left[\begin{matrix}{\mathbf{A}}_{1}{\mathbf{x}}_{1}&{\mathbf{A}}_{1}{\mathbf{x}}_{2}&\cdots&{\mathbf{A}}_{1}{\mathbf{x}}_{K}\\ {\mathbf{A}}_{2}{\mathbf{x}}_{1}&{\mathbf{A}}_{2}{\mathbf{x}}_{2}&\cdots&{\mathbf{A}}_{2}{\mathbf{x}}_{K}\\ \vdots&\vdots&\ddots&\vdots\\ {\mathbf{A}}_{K}{\mathbf{x}}_{1}&{\mathbf{A}}_{K}{\mathbf{x}}_{2}&\cdots&{\mathbf{A}}_{K}{\mathbf{x}}_{K}\\ \end{matrix}\right].$
Finally, the color image is
(19) ${\mathbf{Z}}={\mathbf{W}}{\mathbf{Q}}=\mathrm{diag}\left({\mathbf{A}}{\mathbf{X}}\right){\mathbf{Q}}.$
where ${\mathbf{Z}}=\left[{\mathbf{z}}_{1},{\mathbf{z}}_{2},{\mathbf{z}}_{3}\right]\in\mathbb{R}^{MN\times 3}$.
Appendix B Effect of clipping to non-negative values.
It is necessary to clip negative values in the generated metamer data to ensure the resulting spectra are physically plausible (i.e., no negative spectral radiance). This may lead to slight deviations in the RGB values, and therefore images that are not exact metamers. However, we verify that the resulting difference is actually negligible by comparing the projected RGB images from the metamer pairs. For example, in the experiments of Table 5 in the main paper, 32.9% of the generated metamers produce exactly the same RGB images (exact-metamers). Among the remaining 67.1% that are affected by clipping (i.e., near-metamers), the average PSNR between the RGB pairs is 75.8 dB, with a standard deviation of $\pm$17.7 dB. This indicates that the effect of clipping the negative values in the metamer spectra is negligible.